OpenStudy (anonymous):

Find the length of c if the measure of angle A equals 62 degrees and the measure of angle B equals 39 degrees and b = 15.

OpenStudy (anonymous):

I think I got this one but I'm not too sure..

hartnn (hartnn):

what does angle C equal ?

hartnn (hartnn):

what did you get ?

OpenStudy (anonymous):

23.4?

hartnn (hartnn):

for length of c ?

OpenStudy (anonymous):

Yess

OpenStudy (anonymous):

But I really doubt my calculations.

hartnn (hartnn):

then show us what you did , if there's error, we will help you spot it :)

OpenStudy (anonymous):

Okay! So the previous question was the same problem, but asking for the length of a. So I did... C=180-(62+39)=79 . . . a/sin(62=16sin(39=c/sin(79 . . . a=15sin(62/sin(39 . . . c=15sin(79/sin(39 . . . a=21.04 , c=23.4

OpenStudy (anonymous):

...but I doubt my calculations with that, too. Like I said, this section of Geometry is tougher for me.

hartnn (hartnn):

you're good at this now! :D

OpenStudy (anonymous):

:D Thank you!

OpenStudy (anonymous):

What do you think of this newer problem, though?

hartnn (hartnn):

both values of a and c are correct :)

OpenStudy (anonymous):

Okay, great! You've been a huge help, thank you soooo much! c:

hartnn (hartnn):

always happy to help ^_^